Coal is mostly carbon (C). When it is burned, it reacts with oxygen in the air in a combustion reaction.   43% of the electricity generated uses coal. On average, each person uses 2680 kg of coal per year to generate electricity that they use.

How much CO2 is produced?


1
Expert's answer
2021-06-10T03:34:55-0400

The molar masses are:

C - 12.01 g/mol = 0.01201 kg/mol

CO2 - 44.01 g/mol = 0.04401 kg/mol

The reaction equation is:

C + O2 --> CO2

"m(CO_2)=2680kg(C)\times\frac{1 mol(C)}{0.01201kg(C)}\times\frac{1mol(CO_2)}{1mol(C)}\times\frac{0.04401kg(CO_2)}{1mol(CO_2)}=9820kg"


Answer: 9820 kg